OpenStudy (anonymous):

My brain is tired for some reason I can't quite do this one. 5 sqrt -1/32

mathslover (mathslover):

\[\huge{5\sqrt{\frac{-1}{32}}}\] is this the question ?

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

relax and then try!! :D

OpenStudy (anonymous):

No sorry -1/-32 the negative is on both

mathslover (mathslover):

kkk then here : \[\huge{5*\frac{1}{4}\sqrt{\frac{-1}{2}}}\]

mathslover (mathslover):

\[\huge{\frac{5}{4}*\frac{i}{sqrt{2}}}\] \(\huge{\text{here i is iota }}\)

OpenStudy (anonymous):

(5 sqrt(-1)) / 32 = 0.15625 i

OpenStudy (anonymous):

Thanks fellas you pushed me in the right direction... :)